OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! Find the area under the graph of f(x) = e-lnx(x) on the interval [1, 2]. a. 0.5 b. -0.5 c. 0.693 d. -0.693

OpenStudy (aryana_maria2323):

Can you please help? @zepdrix @ganeshie8 @Directrix

Nnesha (nnesha):

\[area = \int\limits_{a}^{b} f(x) dx\] http://prnt.sc/axbnia

Nnesha (nnesha):

take the anti derivative of the function and then find F(b)-F(a) where F represents the(anti derivative )

OpenStudy (aryana_maria2323):

Okay so will I replace x with 1 and 2?

Nnesha (nnesha):

\[f(x)=e -\ln(x)\] this is the equation right

OpenStudy (greatlife44):

yeah that's what I want to know too lol

Nnesha (nnesha):

i think that is a typo

OpenStudy (greatlife44):

or \[e^{lnx} ?\]

Nnesha (nnesha):

find take the anti derivative and then yes substitute x for 2 and 1

Nnesha (nnesha):

hm

satellite73 (satellite73):

\[e^{-\ln(x)}=e^{\ln(\frac{1}{x})}=...\]

OpenStudy (aryana_maria2323):

\[e^-{lnx(x)}\]

OpenStudy (greatlife44):

\[e^{-\ln(x)} = e^{\ln(\frac{ 1 }{ x })} = \frac{ 1 }{ x } = ? \]

OpenStudy (aryana_maria2323):

e^-lnx(x) this is the equation.

OpenStudy (aryana_maria2323):

I

OpenStudy (greatlife44):

essentially that equals 1/x

OpenStudy (aryana_maria2323):

I am truly confused. Can someone please show me in steps on how to solve this?

zepdrix (zepdrix):

\(\large\rm e^{-(ln x)x}\) This?

zepdrix (zepdrix):

\(\large\rm x e^{-(ln x)}\) Or this maybe?

OpenStudy (greatlife44):

I think this is what the question was \[e^{-\ln(x)} = e^{\ln\frac{ 1 }{ x }} = \frac{ 1 }{ x }\] \[\int\limits_{1}^{2} \frac{ dx }{ x }\]

zepdrix (zepdrix):

Aryana, you need to make your question more clear...

OpenStudy (aryana_maria2323):

The equation is \[e ^{-lnx(x)}\]

OpenStudy (greatlife44):

you can actually just re-write that as 1/x

OpenStudy (aryana_maria2323):

Ohh okay then I guess that is what my new equation would be.

zepdrix (zepdrix):

Both x's are in the exponent like that? Ok.

OpenStudy (greatlife44):

This is what you're integrating \[\int\limits_{2}^{1} \frac{ dx }{ x }\]

OpenStudy (greatlife44):

sorry, bounds of integration are wrong lol 2 should be on top.

zepdrix (zepdrix):

\[\large\rm e^{-\ln(x)\cdot x}\quad=e^{\ln(x^{-1})\cdot x}\quad=e^{\ln(x^{-x})}=x^{-x}\]

zepdrix (zepdrix):

It's certainly not 1/x that you're integrating. Not if you have 2's x's up there like that.

zepdrix (zepdrix):

2* x's*

OpenStudy (aryana_maria2323):

Yes the x's are right next to each other like x(x). That is what was confusing me.

satellite73 (satellite73):

what we need is a screen shot perhaps you can't really have two x's next to each other i mean you can, but you shouldn't

OpenStudy (aryana_maria2323):

One second I will send a screen shot.

OpenStudy (aryana_maria2323):

@satellite73 @zepdrix @greatlife44 @Nnesha

zepdrix (zepdrix):

Personally, I think it's a typo, because if it had been \(\large\rm e^{-ln x}\), then it would actually correspond to one of your options, as greatlife was indicating.

OpenStudy (greatlife44):

so actually if you were to just integrate this

Nnesha (nnesha):

http://prnt.sc/axbvxg

OpenStudy (greatlife44):

\[\int\limits_{1}^{2} \frac{ dx }{ x } = \ln|x|+C\]

OpenStudy (aryana_maria2323):

I got 0.5 but I don't know if that is correct. I think I might have input it wrong.

OpenStudy (aryana_maria2323):

@greatlife44 @Nnesha @zepdrix

OpenStudy (greatlife44):

so first the integral of 1/x is ln|x|+C now we evaluate it

OpenStudy (greatlife44):

\[\ln|x|+C \] from 2 to 1 =\[Ln(2)-Ln(1)\]

OpenStudy (greatlife44):

ln(1) = 0 so you're left with Ln(2). just exactly how did you get 0.5?

OpenStudy (aryana_maria2323):

Ohh okay. I see how you got Ln(2) which equals .0693

satellite73 (satellite73):

it really says \[e^{-x\ln(x)}=x^{-x}\] i have no idea to how to that